HUBBLE SPACE TELESCOPE DAILY REPORT #1034 PERIOD COVERED: 0800 EST Daily Status Report as of 350/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of FOS/S/C 4762 (FOS Pressure Measurement) The Faint Object Spectrograph was placed in the LVON Mode, and the ion gauge was turned on for over 2.5 hours on each side to measure the pressure within the aft shroud. No anomalies were reported. 1.2 Completed FOS/BL 4802 (SMOV FOS/BL Internal Operations) The Faint Object Spectrograph was used to make dark exposures and calibration lamp exposures to verify the science data interface and the proper functioning of the FOS. The dark images were received and displayed. They were reported to look as expected. The wave length calibration images also looked as expected. 1.3 Completed FOS/BL 5243 (FOS Location of Spectra and Aperture: SMOV Tests) The Faint Object Spectrograph was used to do a SMOV internal calibration observation before the COSTAR deployment to determine the course location of the 0.3 arcsecond aperture and the Y-location of four spectral grating settings. A total of five observations were received and displayed. Hard copies were made. The data was reported to look normal. 1.4 Completed S/C 4756 (GHRS SMOV Pressure Test) A pressure test was run on the High Resolution Spectrograph to determine how the pressure has dropped since the aperture and vent doors were opened. The pressure was reported to be 1.0E-06 Torr. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: No Guide Star Acquisitions were taken during this period. 2.2 FHST Updates: No Fixed Head Star Tracker updates were taken during this period. 2.3 Significant HSTARs: None 2.4 Operations Note: The HST Ground System continues to operate nominally. SI operations are being controlled principally by STScI-generated SMSs. Initial high voltage activities for the FOS Blue Side and GHRS side 2 were executed with no operational issues. Data has been forwarded to the STScI OSS for further evaluation. The FHST/RGA scale factor alignment process is producing data that continues to be analyzed. Both gyros and magnetometers are exhibiting nominal performance. The rate biases for gyros 4, 5, and 6 are converging to a stable state. The EPS status is nominal. There were no configuration changes. Average maximum battery SOC is about 500 Ahr, and SA output is approximately 154 amps. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete S/C 5253 (Internal FHST/RGA Scale Factor Alignment Determination) 3.2 Complete FOS/RD 4803 (SMOV FOS/RD Internal Operations) 3.3 Complete FOS/RD 5243 (FOS Location of Spectra and Aperture: SMOV Tests) 3.4 Complete Two Sets of GHRS/2 4780 (SMOV GHRS Detector Calibrations) 3.5 Complete GHRS/2 4757 (GHRS SMOV Light Path Verification) 3.6 Complete FOC/96 5077 (FOC/96 SMOV Reverification Proposal) 3.7 Complete FOC/96 4679 (FOC SMOV Internal Flat Fields) 3.8 Complete S/C 4711 (WFPC2 Decontamination & KSPOT Check) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END- .